7y=2y-15,9x=2x+14,99+6u=-3u
Answer by rahman(247)

7y=2y-15
7y-2y=-15
5y=-15
y=-3
==========
9x=2x+14
9x-2x=14
7x=14
x=2
==========
99+6u=-3u
99=-3u-6u
99=-9u
u=-11